Surface area of a cube is 27 square centimeters. Calculate the length of the sides of the cube
1
Expert's answer
2021-07-30T15:17:57-0400

surface area of a cube =area of one side "\times 6"

let the area of one side be x

"x\times 6=27\\6x=27\\x=4.5cm^2"

"area=length squared\\=l^2\\4.5=l^2\\\sqrt{4.5}=l\\l=2.12"